OpenStudy (anonymous):

The distance between the points (1,8) and (4,4) is equal to 5.

OpenStudy (anonymous):

True or false?

OpenStudy (helder_edwin):

do you know how to compute the distance between 2 points?

OpenStudy (anonymous):

no i dont. can u explain it to me?

OpenStudy (anonymous):

\[\sqrt{(\Delta x)^2+(\Delta y)^2}\]

OpenStudy (anonymous):

woahh

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

k...do you know what to do from here?

OpenStudy (anonymous):

no, ive actually never seen that. im just going to try it

OpenStudy (anonymous):

so, change in x is 4-1=3 change in y is:4-8=-4 square them. add the result. take square root.

OpenStudy (anonymous):

thank you very much. i appreciate it

OpenStudy (anonymous):

no prob; let me know what you get

OpenStudy (anonymous):

okay im doing it now

OpenStudy (anonymous):

i do the -4+3=-1?

OpenStudy (anonymous):

no. square each one individually first.

OpenStudy (anonymous):

\[3^2+(-4)^2\]

OpenStudy (anonymous):

then take the square root of whatever you get

OpenStudy (anonymous):

9+-16=-7

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no wait, i did it wrong because i cant square a neg. number

OpenStudy (anonymous):

9+16 because\[(-4)^2=(-4)(-4)=16\]the square of a negative is a positive number

OpenStudy (anonymous):

oh sorry. i suck at math. so i square the 16 and get 4?

OpenStudy (anonymous):

you square the -4. The gives you 16. the negative sign goes away.

OpenStudy (anonymous):

so now you should have:\[\sqrt{9+16}\]

OpenStudy (anonymous):

9+16=25 25=5

OpenStudy (anonymous):

well I assume you mean\[\sqrt{25}=5\]but yeah, there's your distance

OpenStudy (anonymous):

Yes it´s 5.

OpenStudy (anonymous):

sqrt((x2-x1)^2+(y2-y1)^2)= sqrt ((4-1)^2+(4-8)^2

OpenStudy (anonymous):

sqrt((3)^2 +sqrt(-4)^2) = sqrt((9) +sqrt(16))= sqrt (25) = 5